OpenStudy (anonymous):

Find an exact value. sin 17 pi / 12 this is sum and difference formula for trigonometry

OpenStudy (anonymous):

\[\frac{2}{3}+\frac{3}{4}=\frac{17}{12}\] right?

OpenStudy (anonymous):

and so \[\frac{2\pi}{3}+\frac{3\pi}{4}=\frac{17\pi}{12}\] then find \[\sin(\frac{17\pi}{12})=\sin(\frac{2\pi}{3}+\frac{3\pi}{4})\] using the so called "double angle" formula

OpenStudy (anonymous):

you good from there?